OpenStudy (anonymous):

What is the value of h? 4ʰ=1/64

Nnesha (nnesha):

hey! can you rewrite 64 in terms of base `4`

Nnesha (nnesha):

4 to the what power =64 ?

OpenStudy (anonymous):

4^3 power =64?

Nnesha (nnesha):

right \[\huge\rm 4^h =\frac{ 1 }{ 4^3 }\]we are working on right hand side how move 4^3 to the top of the fraction remember the exponent rule \[\huge\rm x^{-m}=\frac{ 1 }{ x^m }\] when we flip the fraction sign of the `exponent` would change

Nnesha (nnesha):

now* not how

OpenStudy (anonymous):

im confused :/ @Nnesha

OpenStudy (michele_laino):

we have this: \[\frac{1}{{64}} = \frac{1}{{16 \cdot 4}} = \frac{1}{{4 \cdot 4 \cdot 4}} = \frac{1}{{{4^3}}} = {4^{ - 3}}\]

Nnesha (nnesha):

alright i'll give u some examples \[x^{-2} = \frac{ 1 }{ x^{2} }\] it was negative 2 when i flip it it becomes positive 2 \[\large\rm \frac{ 1 }{ x^5 }=\frac{ x^{-5} }{ 1 }\] when i flipped the fraction to move x^5 to the top it becomes x^{-5} so \[\frac{ 1 }{ 4^3 }= ??\]

OpenStudy (anonymous):

okay, thanks could you help me with more? @Nnesha

Nnesha (nnesha):

got it ? hm h= ?

OpenStudy (anonymous):

-3

Nnesha (nnesha):

right.

OpenStudy (anonymous):

What is 4.56 x 10⁸ written in standard form?

Nnesha (nnesha):

alright S.N to Standard form 1st)move the decimal point to the right (the exponents tells us how many places to move right side )

Nnesha (nnesha):

brb

Nnesha (nnesha):

if the exponent is negative moves decimal to the left side and if it's positive then right side

Nnesha (nnesha):

the expo net is positive so we should move the decimal 8 places to the right  now there are only 2 number after decimal for the rest of place we should put 0